OpenStudy (anonymous):

Solving quadratics parabola Y=x^2-4x+6

OpenStudy (solomonzelman):

you want the x-interecepts?

OpenStudy (solomonzelman):

*intercepts

OpenStudy (anonymous):

This is factoring parabolas by the vertex and the zero

OpenStudy (anonymous):

I don't know I just started parabolas an I'm confused

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=x^2-4x+6 }\) to find the x-intercepts, you need to know which value(s) to plug in for x, in order to get 0 (for y). This is why we set y=0, the following way: \(\large\color{black}{ \displaystyle 0=x^2-4x+6 }\) Now, we need to solve this quadratic equation. The fastest method I see is completing the square. (you can choose another methdo if you don't like this one) \(\large\color{black}{ \displaystyle 0=x^2-4x+6 }\) I will make the right hand side into a perfect square trinomial, by subtracting 2 from both sides. \(\large\color{black}{ \displaystyle 0\color{red}{-2} =x^2-4x+6\color{red}{-2} }\) \(\large\color{black}{ \displaystyle -2=x^2-4x+4 }\) then, I will factor the right side \(\large\color{black}{ \displaystyle -2=(x-2)(x-2) }\) \(\large\color{black}{ \displaystyle -2=(x-2)^2 }\)

OpenStudy (solomonzelman):

Now, all that remains is to: \(\small\color{black}{ \bullet }\) Square root both sides. \(\small\color{black}{ \bullet }\) Add 2 to both sides. There are your solutions \(\small\color{black}{ {\bf!} }\)

OpenStudy (solomonzelman):

They all are imaginary however. We can prove that: \(\large\color{black}{ \displaystyle y=x^2-4x+6}\) \(\large\color{black}{ \displaystyle y=x^2-4x+4+2}\) \(\large\color{black}{ \displaystyle y=(x-2)^2+2}\) \(\large\color{black}{ \displaystyle {\rm Vertex}=(2,2)}\)

OpenStudy (solomonzelman):

Since vertex is above the x-axis, and this parabola is opening up (saying that opening up, implies that its Y well get only larger than the vertex but not smaller than the vertex) AND THEREFORE, this parabola does not intersect the x-axis.